Let f (x) =integral
upper limit is x
lower limit is 0
(2u+1/u^2-2u+2)du then find the least value of f (x) in [-1,1].

Grade:Upto college level

1 Answers

Sher Mohammad IIT Delhi
10 years ago
2u+-2+2+1=2(u-1)+3
u^2-2u+2=(u-1)^2+1
2u+1/u^2-2u+2



f'(x)=0, imply x=-1/2


Sher Mohammad
B.Tech, IIT Delhi